It takes bob 4 hours to paint a house and it takes jim 6 hours to paint the same house. How long will it take bob and jim together to paint the house?
Answer by nerdybill(7384)

It takes bob 4 hours to paint a house and it takes jim 6 hours to paint the same house. How long will it take bob and jim together to paint the house?
.
Let x = time (hours) it takes for both
then
x(1/4 + 1/6) = 1
multiplying both sides by 12:
x(3 + 2) = 12
x(5) = 12
x = 12/5
x = 2 and 2/5 hours
or
x = 2.4 hours
or
x = 2 hours and 24 minutes